# Learning Chemical Reaction Representation with Reactant-Product Alignment

### v1

@@ 5.3. Reaction Selectivity Prediction
Furthermore, when dealing with extremely small-scale datasets, the performance of our model can be further enhanced by incorporating more rule-based features, such as the electronic distribution of atoms, as our model does not impose restrictions on the implementation of the MPNN network used to encode reactants and products. Large-scale pretraining on a reaction dataset will also be beneficial. However, these improvement methods are beyond the scope of this paper, and we will reserve them for future work.
5.4. Ablation Study
@@
We investigate the effects of different components in our proposed pipelines. We remove or substitute the distinct components of our model and subject them to testing on the USPTO_500MT dataset and the random split setting of the Buchwald-Hartwig Dataset. The result is summarized in Table 8.
@@
Atom Aligned Encoder. We remove the information fusion layers of Atom Aligned Encoder. Under this circumstance, the Atom Aligned Encoder has been simplified into a network composed of two separate MPNN networks that encode reactants and products, respectively. As observed in Table 8, the removal of the information fusion layer has led to a decrease in the model’s performance across all metrics for different tasks, especially in terms of the top-1 and top-3 accuracy for reaction condition prediction on the USPTO_500MT dataset. This suggests that by incorporating the alignment of atoms before and after the reaction into the model, the Atom Aligned Encoder can more effectively discern the differences between the molecules before and after the reaction, thus offering more robust reaction representations for downstream tasks.
@@
Reaction-Center-Aware Decoders. We replace the RC-aware cross-attention layers with the original cross-attention layer proposed in (Vaswani et al., 2017). Table 8 demonstrates a decline in model performance in terms of all metrics for both the sequential generation tasks and the task that requires a reaction-level representation. This clearly demonstrates that the RC-aware cross-attention mechanism enables the model to focus on the core functional groups of the reaction and comprehend the reaction process, thereby leading to performance improvements.
6. Conclusion
@@
In this paper, we propose RAlign, a novel chemical reaction representation learning model. Our model integrates the atomic correspondence between reactants and products, as well as information about the reaction center, enabling the model to better model the reaction process and gain a deeper understanding of the reaction mechanism. An adapter is utilized to incorporate reaction conditions, allowing our model to adapt to various modalities of reaction conditions and efficiently leverage previous work to enhance performance. Experimental results demonstrate that our model architecture outperforms existing reaction representation learning architectures across various downstream tasks. In the future, we plan to use this architecture for large-scale pretraining to use the reaction representation space to assist scientists in conducting research on reaction mechanisms.
@@
Limitations. The model requires atom mappings as input. Although there are now tools for accurate atom-mapping, incorrect atom-mapping can still have a negative impact on the model’s performance. Like most deep-learning methods, our model requires a substantial amount of training data for support; hence, on small-scale datasets, our model still cannot surpass hand-crafted features.
Appendix A Statistical Information of Datasets
@@
We summarize the statistical information of all the dataset set used in this work in Table 9.
@@
Appendix B Implementation Details
B.1. Pipeline for different tasks
This work does not train a multi-task unified model like T5chem (Lu and Zhang, 2022). Given the different datasets and varying inputs and outputs for each task, we have selected different modules to assemble the pipelines for the tasks discussed in this paper. The two pipelines involved are displayed in Fig. 2. The input for the reaction condition prediction task consists solely of reactants and products, hence the pipeline for this task is composed of an atom-aligned encoder followed by a RC-aware decoder with sequential output. For the reaction yield prediction and reaction selectivity prediction tasks, which also include reaction conditions in their input, we have incorporated a reaction condition encoder and utilized a RC-aware decoder with single output.
B.2. Initial node/edge feature extraction
We use the atom and bond encoder provided by Open Graph Benchmark (Hu et al., 2020a). These encoders transform the atom and chemical bonds into integer numbers based on their characteristics and molecular structure. Nine atom descriptors are provided, including the atom type, formal charge and other properties that can be calculated by RDKit (Landrum et al., 2013). As for chemical bonds, three descriptors are used, including bond type, bond stereochemistry as well as whether the bond is conjugated. Every descriptor corresponds to a learnable embedding table. The initial node and edge features are derived by aggregating the embeddings corresponding to each descriptor.
B.3. MPNN Layer in Atom Aligned Encoder
@@
B.4. Condition Encoders for Buchwald–Hartwig and thiol addition Selectivity Dataset
The Buchwald–Hartwig and thiol addition selectivity dataset uses chemical reagents as reaction conditions. To generate condition features, we employ a graph neural network. It should be noted that the number of distinct reagents in these datasets is quite limited, with less than 50 different molecules available, which makes it challenging to train a reaction condition encoder from scratch. Therefore, we have opted to use a pretrained molecular representation model proposed by Hu et al.. This model is lightweight and has been pretrained using only the random atom masking task. The roll-out form of the -th layer of this model is formulated as
@@ B.4. Condition Encoders for Buchwald–Hartwig and thiol addition Selectivity Dataset
B.5. Model Implementation details
We implement our model based on torch_geometric 2.2.0 (Fey and Lenssen, 2019) and Pytorch 1.13 (Paszke et al., 2019). For model for reaction condition prediciton, we set the hidden size as 512, encoder layer as 6, decoder layer as 6 and the number of attention heads as 8. The dropout ratio is set as 0.1. The highest learning rate of each model as set as 1.25e-4. We slowly increase our learning rate to the highest in the first few epochs and slowly decrease it using exponential decay. For reaction yield prediction, we set the hidden size as 128, encoder layers as 3 and attention heads as 8. The dropout ratio is set as 0.1 and the learning rate is set as 1e-4. For reaction selectivity prediction in Thiol addition dataset, we set the hidden size as 128, encoder layers as 3 and attention heads as 8. The dropout ratio is set as 0.0 and the learning rate is set as 5e-5. For reaction selectivity prediction in C-H functionalization dataset, we set the hidden size as 128, encoder layers as 5 and attention heads as 8. The dropout ratio is set as 0.0 and the learning rate is set as 5e-4. All models are trained with Adam optimizer (Kingma and Ba, 2014). We will consider open-sourcing our code upon the paper acceptance.
Appendix C Data Preparation
C.1. Adding Atom-mapping
The Buchwald-Hartwig dataset features a consistent reaction template across all its reaction data, enabling the derivation of atom-mapping through rule-based approaches. For all the other datasets except the Buchwald-Hartwig dataset, we have employed Rxnmapper (Schwaller et al., 2021a) to obtain atom-mapping. It is noteworthy that we have re-annotated the labels for the USPTO_500MT dataset, treating both the originally provided reactants and reagents as input reactants during the atom mapping annotation for this dataset. For further details on the processing of the USPTO_500MT dataset, please refer to Appendix C.2.
C.2. Generation of Reagents of USPTO_500MT
@@
Appendix D Further elaboration on reaction condition prediction
To provide a more detailed exposition of the input, output, and evaluation metrics for the reaction condition prediction task, we present several examples in Fig. 3 and Fig. 4. In the USPTO_CONDITION dataset, a reaction condition is defined by the presence of a catalyst, reagent, and solvent. A reaction condition combination is deemed correct if all components are accurately predicted, with a component being considered correctly predicted only if every molecule within it is predicted without error. Conversely, for the USPTO_500MT dataset, which does not classify reaction conditions further, a prediction is deemed correct if the predicted molecules match exactly with those in the labeled data.

### v2

@@ 5.3 Reaction Selectivity Prediction
Furthermore, when dealing with extremely small-scale datasets, the performance of our model can be further enhanced by incorporating more rule-based features, such as the electronic distribution of atoms, since our model does not impose restrictions on the implementation of the MPNN network used to encode reactants and products. Large-scale pretraining on a reaction dataset will also be beneficial. However, these improvement methods are beyond the scope of this paper and will be reserved for future work.
5.4 Ablation Study
@@
We investigate the effects of different components in our proposed pipelines. We remove or substitute the distinct components of our model and subject them to testing on the USPTO_500MT dataset and the random split setting of the Buchwald-Hartwig dataset. The results are summarized in Table 8.
@@
Atom Aligned Encoder. We remove the information fusion layers of the Atom Aligned Encoder. Under this circumstance, the Atom Aligned Encoder has been simplified into a network composed of two separate MPNN networks that encode reactants and products, respectively. As observed in Table 8, the removal of the information fusion layer has led to a decrease in model performance in both tasks, especially in terms of the top-1 and top-3 accuracy for reaction condition prediction on the USPTO_500MT dataset. This suggests that by incorporating the alignment of atoms before and after the reaction into the model, the Atom Aligned Encoder can more effectively discern the differences between the molecules before and after the reaction, thus offering more robust reaction representations for downstream tasks.
@@
Reaction-Center-Aware Decoders. We replace the RC-aware cross-attention layers with the original cross-attention layer proposed in (Vaswani et al., 2017). Table 8 demonstrates a decline in model performance in terms of all metrics for both sequential generation tasks and tasks that require a reaction-level representation. This clearly demonstrates that the RC-aware cross-attention mechanism enables the model to focus on the core functional groups of the reaction and to comprehend the reaction process, thereby leading to performance improvements.
5.5 Case Study
To intuitively demonstrate what our proposed module has learned, we conducted visualizations of both the embeddings learned by the atom-aligned encoder and the attention coefficients from the RC-aware cross-attention mechanism.
The Atom-Aligned encoder helps the model learns beyond the topological structures of molecules. To explore the changes brought about by atomic alignment for the encoder, we visualized the node embeddings generated by the atom-aligned embedding and conventional MPNN networks. The models for feature extraction were trained on the reaction condition generation task for the ablation study discussed in Sec. 5.4, and the visualizations are displayed in Fig. 2. From the figure, we can observe that due to MPNN networks only receiving topological structure information from individual molecules and the high similarity in topology between reactants and products, the node embeddings of the MPNN network are updated along almost the same pattern trajectory. Especially in the 3rd to 6th layers of the network, the spatial structure formed by the nodes does not undergo significant changes. In contrast, the atom-aligned encoder facilitates the exchange of information between different components in chemical reactions, leading to a more diverse feature space and thereby enhancing the model’s representational capacity.
RC-aware Cross-attention helps the model to find key functional groups. We visualize the average attention weights of different kinds of heads for predicting the catalyst of the given reaction in Fig. 3. The presented case involves the reduction of aromatic nitro groups to amino groups using catalytic hydrogenation, a method commonly used in industry and relatively eco-friendly and efficient (Orlandi et al., 2016). However, this method is not suitable for situations where the benzene ring has other competitive groups, such as halogens (Shi, 2007; Yao & Zhang, 1984). The visualization reveals that while our RC-aware heads are restricted to focus on the reaction center, thereby providing the model with information about the reaction type, our normal heads attend to other groups on the benzene ring. This distribution of attention weights allows the model to correctly confirm the absence of other competitive groups on the benzene ring, thereby accurately predicting that the reaction can proceed via catalytic hydrogenation and precisely predicting the corresponding catalyst. The aforementioned case demonstrates that RC-aware cross-attention enables different attention mechanisms to assign clear roles to different attention heads, thereby helping the model detect and focus on key functional groups beyond the reaction center. We also provide the visualization of attention heads for predicting other components of the reaction condition and visualization for other cases in Appendix A.1.
6 Conclusion
@@
In this paper, we propose RAlign, a novel chemical reaction representation learning model. Our model integrates the atomic correspondence between reactants and products, as well as information about the reaction center, enabling the model to better model the chemical bond changes and gain a deeper understanding of the reaction mechanism. An adapter is utilized to incorporate reaction conditions, allowing our model to adapt to various modalities of reaction conditions and to efficiently leverage previous work to enhance performance. Experimental results demonstrate that our model architecture outperforms existing reaction representation learning architectures across various downstream tasks. In the future, we plan to use this architecture for large-scale pretraining to use the reaction representation space to assist scientists in conducting research on reaction mechanisms.
@@
Limitations. The model requires atom mappings as input. Although there are now tools for accurate atom mapping, incorrect atom mapping can still have a negative impact on the model’s performance. Like most deep learning methods, our model requires a substantial amount of training data for support; hence, on small-scale datasets, our model cannot surpass handcrafted features.
Impact Statement
This paper introduces a novel chemical reaction representation learning architecture, designed to enhance the performance of tasks that utilize chemical reactions as input, such as chemical reaction condition prediction. These tasks are intrinsically linked to the design of molecular synthesis schemes and the optimization of production processes in the organic chemical industry. Our work has the potential to increase the production efficiency of the organic synthesis industry and enable the training of more powerful reward models that align with specific demands of industrial production, thereby feeding back into upstream molecular design and related fields. It is important to note that, at present, our research does not directly engage with broader societal issues, and to date, no adverse social implications have been associated with our work.
Appendix A Supplementary Experimental Results
A.1 Visualization of RC-aware Cross-attention
We have supplemented the cross-attention coefficients for predicting the solvent and reagent for the case presented in Sec. 5.5, as depicted in Fig. 4. The identification of the catalyst and the provision of the reaction center have made the model aware that this reaction can be achieved through catalytic hydrogenation. Consequently, when predicting the solvent, the model’s normal attention heads focus on areas other than the benzene ring, ensuring that the proposed solvent improves reaction selectivity (i.e., the reaction proceeds at the correct site) and provides an appropriate system for its progress. This can also be observed in the normal head’s prediction of the reagent. Additionally, it is noticeable that when predicting the reagent, the model’s RC-aware heads allocate more attention to the reactant, thereby accurately predicting the reagent, hydrogen, which directly reacts with the reactant. We also provide another two cases in Fig. 5 and Fig. 6.
A.2 Embedding Visualization of Atom Aligned Encoder
We have completed the reaction information for the node embeddings shown in Sec. 5.5 and additionally provided visualizations for two more cases, which are presented in Fig. 7. We can observe that in these cases, the point features of the conventional MPNN network still follow the same patterned trajectory for updates. In contrast, our feature space appears more diverse, which is probably why our model is able to achieve better performance.
A.3 Robustness of Our Model against Incorrect Atom-mapping
Existing atom-mapping tools are not entirely reliable, as there may be instances where these tools incorrectly match similar substructures between reactants and products. To investigate the robustness of our model against erroneous inputs from these atom-mapping tools, we explored the relationship between the accuracy of the model in predicting/generating chemical reaction conditions and the atom mapping confidence scores provided by RXNMapper (Schwaller et al., 2021a). We segmented the data from USPTO_CONDITION and USPTO_500MT based on the atom mapping confidence scores in increments of 0.1 and then calculated the top- accuracy within each interval. The results are depicted in Fig. 8. Additionally, we computed the Pearson correlation coefficients between the median confidence and the top-k accuracy of samples within each interval, which are presented in Table 9.
| Dataset | top- | top- | top- | top- | |
| USPTO_CONDITION | 0.096 | 0.053 | -0.020 | -0.077 | |
| USPTO_500MT | -0.037 | 0.283 | 0.367 | 0.415 |
From the table and figure, we observe that there is no strong correlation between our model’s performance and the confidence of atom mapping. This indicates that although our model relies on atom mapping as input, it maintains a certain level of robustness against erroneous atom mapping, as long as the atom mapping can match similar substructures between reactants and products.
Appendix B Datasets
B.1 Statistical Information of Datasets
@@
We summarize the statistical information of all the datasets used in this work in Table 10.
@@
B.2 Metrics for evaluation
We use various metrics to evaluate task performance on different datasets. The USPTO_CONDITION dataset serves as the benchmark for evaluating the Condition Prediction task, the USPTO_500MT dataset is used for the Condition Generation task, the Buchwald-Hartwig dataset is employed for the Reaction Yield Prediction task, and the Thiol Addition and C-H Functionalization datasets are used for the Reaction Selectivity Prediction task.
Condition Prediction/Generation. We assess the predictive performance for the whole reaction condition combinations as well as for each constituent element within the reaction conditions. We use the conventional top- accuracy to evaluate the performance. A prediction is considered correct if and only if all of the molecules of it are correctly predicted. The order of the constituents is not one of the criteria for determining whether a prediction is correct. When a reaction in the test set has multiple recorded reaction combinations, a prediction is considered correct if it is completely consistent with any one of them.
Reaction Yield/Selectivity Prediction. Datasets for these tasks provide the single numeric reaction output (yield or selectivity metric) as prediction targets, thus we could consider these tasks as regression problems, and use the conventional metrics including score, root mean square error (RMSE), and mean absolute error (MAE), for evaluation.
Appendix C Implementation Details
C.1 Pipeline for different tasks
This work does not train a multi-task unified model like T5chem (Lu & Zhang, 2022). Given the different datasets and varying inputs and outputs for each task, we have selected different modules to assemble the pipelines for the tasks discussed in this paper. The two pipelines involved are displayed in Fig. 9. The input for the reaction condition prediction task consists solely of reactants and products, hence the pipeline for this task is composed of an atom-aligned encoder followed by a RC-aware decoder with sequential output. For the reaction yield prediction and reaction selectivity prediction tasks, which also include reaction conditions in their input, we have incorporated a reaction condition encoder and utilized a RC-aware decoder with single output.
C.2 Initial node/edge feature extraction
We use the atom and bond encoder provided by Open Graph Benchmark (Hu et al., 2020a). These encoders transform the atom and chemical bonds into integer numbers based on their characteristics and molecular structure. Nine atom descriptors are provided, including the atom type, formal charge, and other properties that can be calculated by RDKit (Landrum et al., 2013). As for chemical bonds, three descriptors are used, including bond type, bond stereochemistry, as well as whether the bond is conjugated. Every descriptor corresponds to a learnable embedding table. The initial node and edge features are derived by aggregating the embeddings corresponding to each descriptor.
C.3 MPNN Layer in Atom Aligned Encoder
@@
C.4 Condition Encoders for Buchwald–Hartwig and thiol addition Selectivity Dataset
The Buchwald–Hartwig and thiol addition selectivity dataset uses chemical reagents as reaction conditions. To generate condition features, we employ a graph neural network. It should be noted that the number of distinct reagents in these datasets is quite limited, with fewer than 50 different molecules available, which makes it challenging to train a reaction condition encoder from scratch. Therefore, we have opted to use a pretrained molecular representation model proposed by Hu et al.. This model is lightweight and has been pretrained using only the random atom masking task. The roll-out form of the -th layer of this model is formulated as
@@ C.4 Condition Encoders for Buchwald–Hartwig and thiol addition Selectivity Dataset
C.5 Model Implementation details
We implement our model based on torch_geometric 2.2.0 (Fey & Lenssen, 2019) and Pytorch 1.13 (Paszke et al., 2019). For the model for reaction condition prediction, we set the hidden size as 512, encoder layers as 6, decoder layers as 6, and the number of attention heads as 8. The dropout ratio is set as 0.1. The highest learning rate of each model is set as 1.25e-4. We slowly increase our learning rate to the highest in the first few epochs and slowly decrease it using exponential decay. For reaction yield prediction, we set the hidden size as 128, encoder layers as 3, and attention heads as 8. The dropout ratio is set as 0.1, and the learning rate is set as 1e-4. For reaction selectivity prediction in the Thiol addition dataset, we set the hidden size as 128, encoder layers as 3, and attention heads as 8. The dropout ratio is set as 0.0, and the learning rate is set as 5e-5. For reaction selectivity prediction on the C-H functionalization dataset, we set the hidden size as 128, encoder layers as 5, and attention heads as 8. The dropout ratio is set as 0.0 and the learning rate is set as 5e-4. All models are trained with the Adam optimizer (Kingma & Ba, 2014). We will consider open-sourcing our code upon the paper acceptance.
Appendix D Data Preparation
D.1 Adding Atom-mapping
The Buchwald-Hartwig dataset features a consistent reaction template across all its reaction data, enabling the derivation of atom-mapping through rule-based approaches. For all the other datasets except the Buchwald-Hartwig dataset, we have employed Rxnmapper (Schwaller et al., 2021a) to obtain atom-mapping. It is noteworthy that we have re-annotated the labels for the USPTO_500MT dataset, treating both the originally provided reactants and reagents as input reactants during the atom mapping annotation for this dataset. For further details on the processing of the USPTO_500MT dataset, please refer to Appendix D.2.
D.2 Generation of Reagents of USPTO_500MT
@@
Appendix E Baselines
E.1 Chemical Reaction Condition Prediction
We choose T5Chem (Lu & Zhang, 2022), Parrot-LM_E (Wang et al., 2023), GCNN (Maser et al., 2021), Reagent Transformer (Andronov et al., 2023), and FPRCR (Gao et al., 2018) as our baselines for Chemical Reaction Condition Prediction; the details of them are as follows.
Parrot-LM_E is a transformer-based model. It takes the SMILES of the reaction as input to predict the chemical reaction condition combinations. It is specifically designed for reaction condition combinations with a fixed number of components and it is developed from the checkpoint of BERT (Devlin et al., 2019).
Reagent Transformer is also a transformer-based model. It is developed from the checkpoint provided by Molecular Transformer (Schwaller et al., 2019), a model pretrained on predicting chemical reaction products. It should be noted that in the original design of Reagent Transformer, the decoder is used to generate SMILES strings. For a fair comparison, we replaced the original vocabulary optimized for SMILES generation with a predefined molecular library and adapted the reaction condition prediction problem into a sequence generation task with a fixed length of 5.
T5Chem is a language model pretrained on the PubMed dataset (Kim et al., 2020) with a self-supervised task and the USPTO_500MT dataset (Lu & Zhang, 2022) with five different supervised tasks including chemical reaction condition generation. Similar with Reagent Transformer, the decoder of T5Chem is designed for SMILES generation. Thus, we replaced the original vocabulary optimized for SMILES generation with a predefined molecular library and adapted the reaction condition prediction problem into a sequence generation task with a fixed length of 5.
FPRCR is a fingerprint-based method specifically designed for reaction condition combinations with a fixed number of components. The model comprises a set of MLPs that correspond to the number of components in the reaction condition combination. All MLPs are invoked in the order of prediction, taking the fingerprints of reactants, products, and the already predicted reaction condition combination components as input to forecast the next molecule that should be used as a condition.
GCNN is a graph-based method that takes the molecular graphs of reactants and products as input to generate chemical reaction representations for predicting reaction condition combinations. The reaction condition combinations in the USPTO_CONDITION dataset are composed of up to one catalyst, two solvents, and two reagents. We have extended the model’s prediction head from a single multilabel classification head to five separate single-label classification heads.
E.2 Chemical Reaction Condition Generation
We choose GCNN (Maser et al., 2021), Reagent Transformer (Andronov et al., 2023), and T5Chem (Lu & Zhang, 2022) for Chemical Reaction Generation Prediction. The details of them are as follows.
Reagent Transformer is a transformer-based model. It is developed from the checkpoint provided by Molecular Transformer (Schwaller et al., 2019), a model pretrained on predicting chemical reaction products. It’s originally desigined for SMILES generation so we do not make any modification.
T5Chem is a language model pretrained on the PubMed dataset (Kim et al., 2020) with a self-supervised task and the USPTO_500MT dataset (Lu & Zhang, 2022) with five different supervised tasks including chemical reaction condition generation. In this paper, we present the results of two versions of T5Chem. One is developed from the pretrained checkpoint and the other is trained from scratch on only the chemical reaction condition generation task.
GCNN is a graph-based method that takes the molecular graphs of reactants and products as input to generate chemical reaction representations for predicting reaction condition combinations. The original design of it is not suitable for SMILES generation, so we follow the original design of it and formulate the chemical reaction condition generation task on USPTO_500MT as multilabel classification problem for it. The labels here is to present whether a molecule in the predefined molecular library could be a part of the reaction condition combination, and use the combinations with the -highest probabilities to calculate top- accuracy.
E.3 Chemical Reaction Yield/Selectivity Prediction
The tasks of predicting chemical reaction yield and selectivity are highly analogous regression tasks; hence, their baselines are identical. It is noteworthy that the emphasis of this paper is on introducing a novel architecture for chemical reaction representation learning, rather than being limited to the specific task of Chemical Reaction Yield/Selectivity Prediction. Therefore, we have not included baselines with complex pretraining tasks or intricate training loss and inference strategies. Pretraining and specialized loss functions tailored for Reaction Yield Prediction are orthogonal to our research. The details of the chosen baselines are as follows.
DRFP (Probst et al., 2022) is a kind of chemical reaction fingerprint for yield prediction, which calculates the fingerprint of the symmetric difference between the n-grams of reactants and products.
MFF (Sandfort et al., 2020) uses multiple fingerprint features of all molecules in one reaction as the input of the regressor.
Chemprop (Heid et al., 2024) is a kind of Message Passing Network encoding reactions. It converts the reactant and product SMILES pairs into a single pseudo-molecule, known as the condensed graph representation (CGR) of the reaction, and processes it through a directed messaging passing network (D-MPNN) to extract reaction features.
RXNFP (Schwaller et al., 2021b) is a transformer model that has been pre-trained on the Pistachio dataset (Mayfield et al., 2017) utilizing self-supervised learning tasks. Building upon this work, the original research team has also proposed a finetuned version for predicting chemical reaction yields, termed YieldBert (Schwaller et al., 2021c).